User Tools

Site Tools


handover:aov027

AOV027

Hobart 26m

Disk VSN: HOB +0115/16000 | Data volume at beginning: 2.6 GB

  1. 17:30 UT Experiment started OK (Atifur)
  2. 17:50 UT He supply pressure seems to be gradually falling (currently at 272 psi). However, temp. is below 70K (turbo pump had to be run before the beginning of the experiment) and pressure at 2.4 micHg. (Atifur)
  3. 23:30 UT disk HOB+0115 with Ke LBA and r1859 erased and aov027 start recorded. switched to HOB+0098 (Warren)

Katherine 12m

Disk VSN: HOB +0120/32000 | Data volume at beginning: 1.3 GB

  1. 17:30 UT Experiment started OK (Atifur)
  2. 00:00 UT HOB +0120 with aum008 erased and aov027 start recorded. switched to HOB+0094 (Warren)
  3. 00:35 UT reboot dbbc and mk5, fmset (w)
  4. 00:49 UT resume schedule with scan 262-0049 (w)

Yarragadee 12m

Disk VSN: HOB +0123/32000 | Data volume at beginning: 1.3 GB

  1. 17:30 UT Experiment started OK (Atifur)
  2. 23:45 UT HOB +0123 with aum008 erased and aov027 start recorded. switched to HOB+0095 (Warren)
You could leave a comment if you were logged in.
/home/www/auscope/opswiki/data/pages/handover/aov027.txt · Last modified: 2018/09/19 01:19 by Warren Hankey